### v1

# Generative Topological Networks

Abstract
Generative models have seen significant advancements in recent years, yet often remain challenging and costly to train and use. We introduce Generative Topological Networks (GTNs) – a new class of generative models that addresses these shortcomings. GTNs are trained deterministically using a simple supervised learning approach grounded in topology theory. GTNs are fast to train, and require only a single forward pass in a standard feedforward neural network to generate samples. We demonstrate the strengths of GTNs in several datasets, including MNIST, celebA and the Hands and Palm Images dataset. Finally, the theory behind GTNs offers insights into how to train generative models for improved performance. Code will be published upon acceptance.
1 Introduction
Figure 1: Samples generated from a standard normal distribution using a GTN trained on CelebA $64\times 64$ with a latent dimension of $100$.
Figure 2: Illustration of the mapping produced by $h$ and of the labeling process for training its approximation $\hat{h}$, where $Y$ is normally distributed and $X$ is uniformly distributed. A point $y$ from the normal sample is labeled with the unique point $x_{y}$ from the uniform sample that has the same empirical CDF value as $y$.
Deep generative models such as Generative Adversarial Networks (GANs) (Goodfellow et al. 2020), Variational Autoencoders (VAEs) (Kingma and Welling 2013), Energy-Based Models (EBMs) (LeCun et al. 2006; Ngiam et al. 2011), autoregressive models (Bengio and Bengio 2000; Larochelle and Murray 2011), normalizing flows (NFs) (Rezende and Mohamed 2015), and diffusion models (Sohl-Dickstein et al. 2015a), have demonstrated remarkable capabilities for generating samples based on training data distributions (Kang et al. 2023; Ho, Jain, and Abbeel 2020; Ramesh et al. 2021; Ho et al. 2019; Kingma and Dhariwal 2018; Kingma et al. 2016; Reed et al. 2017; Van Den Oord, Vinyals et al. 2017; Ramesh et al. 2021).
Diffusion models, in particular, have seen great advancements over the last few years in a variety of settings (Ho, Jain, and Abbeel 2020; Luo et al. 2023; Nichol and Dhariwal 2021; Dhariwal and Nichol 2021; Ho et al. 2022; Rombach et al. 2022), and possess several advantages compared to other methods. For example, compared to GANs and VAEs diffusion models have more training stability as they do not suffer from mode collapse like GANs (Zhang, Li, and Yu 2018), or posterior collapse like VAEs (Lucas et al. 2019).
However, diffusion models have several downsides. Most prominent is their inferior time-efficiency during training and sampling when compared to VAEs and GANs. This is because they employ a gradual and stochastic transition from a Gaussian distribution to the data distribution, performed over multiple time-steps known as the reverse diffusion process. The complexity of this process also makes it challenging to produce interpolations – desirable in the context of continuous data like audio and video. Additionally, diffusion-based methods often generate samples that fall outside the data distribution (e.g. the horizontally-flipped $3$s in (Sohl-Dickstein et al. 2015b) Figure App.1.). Although recent advancements have made significant steps towards enhancing diffusion models, the complex diffusion process is still a key challenging component (Manduchi et al. 2024).
In this work we introduce a particularly efficient, stable and simple class of generative models – Generative Topological Networks (GTNs). GTNs learn a continuous and invertible mapping between two distributions in a supervised manner that facilitates efficient training and sampling. Specifically, given a training set of samples (e.g. images) and a tractable source distribution (e.g. Gaussian), GTNs learn a function $\hat{h}$ such that, given a $y$ sampled from the source distribution, $\hat{h}(y)$ is a sample representing the training data distribution. GTNs are reminiscent of NFs, which aim to transform one distribution into another using a sequence of invertible and differentiable maps. GTNs, however, are constructed entirely differently, and offer many advantages from both practical and theoretical perspectives.
From a practical perspective, GTNs are extremely simple and efficient, requiring a single, vanilla, feedforward architecture trained using supervised learning. This means that GTNs avoid the complexities of introducing stochasticity during training which is present in VAE and diffusion; circumvent the intricacies of training more complex architectures such as those employed by diffusion and GANs; are not a likelihood-based approach, as are VAE, diffusion and NFs; and do not pose constraints on the structure of the neural network, as do NFs. The supervised manner in which GTNs are trained also forms a strong inductive bias that speeds up convergence. GTNs also operate on latent representations of the data, providing even more enhanced training and sampling efficiency. These advantages manifest in our experiments – realistic samples are obtained at early training epochs, and generating a single image takes around $4$ milliseconds on a standard PC and $0.003$ milliseconds on a T4 GPU.
From a theoretical perspective, GTNs provide guarantees and properties that are desirable in the context of generative models. These include: learnability (via the universal approximation theorem), continuity (for continuous interpolations – Figure 5), bijectivity (for diversity and coverage of the data distribution – Figure 3) and properties that serve as guiding topological principles on how to train generative models for better performance (see the Method section and the swiss-roll example).
The remainder of the paper is structured as follows: In the Method section we first develop the theory behind GTNs for the $1$-dimensional case and use it to generate samples from the swiss-roll. We then extend this theory to higher dimensions and use it to generate samples from the multivariate uniform distribution. In the Experiments section we apply GTNs to real datasets, also including demonstrations of continuous interpolation. We follow with the Related Work section and conclude with the Discussion section.
2 Method
We first develop the method in the simple case of a $1$-dimensional space, and proceed to generalize to higher dimensions.
2.1 1-Dimension
Consider a continuous random variable $X$ with values in $\mathbb{R}$. We wish to generate samples from $X$ without knowing its distribution. One solution would be to sample from a known and tractable distribution, such as the standard normal distribution, and then to apply a function that maps this sample to a corresponding sample from $X$. Diffusion models attempt to approximate such a mapping through gradual stochastic manipulations of the standard normal sample back to a sample from $X$. We will show that, under certain general conditions, such a mapping can be explicitly defined, providing a simple deterministic function which we will denote as $h$ (and which we illustrate in Figure 2). We will show that $h$ is in fact a homeomorphism – it is continuous, invertible and has a continuous inverse (see Definition 2.1). This has significant implications, as we will soon explain.
Defining $h$.
In this section we will define a function $h$ that transforms one distribution into another (illustrated in Figure 2). We will also prove that $h$, under fairly general conditions, possesses certain properties that are desirable in the context of generative models by proving that it is a homeomorphism (Theorem 2.1). For example, we will see that $h$ is bijective, so that each sample $y$ is mapped precisely to one sample $x_{y}=h(y)$, generating different samples for different $y$, and guaranteeing that each sample $x$ has a sample $y$ that generates it. We will also discuss other useful consequences of $h$ being a homeomorphism. We begin by defining a homeomorphism in our context and proceed to define $h$ in Theorem 2.1.
Definition 2.1
(Homeomorphism for $\mathbb{R}^{n}$) Let $S,T$ be two subsets of $\mathbb{R}^{n}$. A function $h:S\to T$ is a homeomorphism if: (1) $h$ is continuous; (2) $h$ is bijective; (3) $h^{-1}$ is continuous. When such an $h$ exists, then $S$ and $T$ are called homeomorphic.
Theorem 2.1
Let $X$ and $Y$ be random variables that have continuous probability density functions (pdfs) $f_{X}$, $f_{Y}$ and supports $S_{X}$, $S_{Y}$ that are open intervals in $\mathbb{R}$. Denote the corresponding cumulative distribution functions (CDFs) as $F_{X}$ and $F_{Y}$. Define:
$\begin{split}h:S_{Y}&\to S_{X}\\ h(y)&={F_{X}}|_{S_{X}}^{-1}({F_{Y}}|_{S_{Y}}(y))\end{split}$ (1)
Then: (1) $h$ is well-defined; (2) $h$ is a homeomorphism.
Note that the requirement that $S_{X},S_{Y}$ are open intervals can be adapted to a union of open intervals – see Appendix A.3. The proof of Theorem 2.1 is in Appendix A.1.
The simple special case of $f_{X},f_{Y}>0$ in $\mathbb{R}$ illustrates the key ideas of Theorem 2.1. In this special case we have that:
$\begin{split}h:\mathbb{R}&\to\mathbb{R}\\ h(y)&=F_{X}^{-1}(F_{Y}(y))\end{split}$ (2)
For simplicity of exposition, we continue with this special case where $f_{X},f_{Y}>0$ on all of $\mathbb{R}$, but note that adapting the results to the general case is a matter of technicality. Namely – $f_{X},f_{Y}>0$ on $S_{X},S_{Y}$ respectively. Combined with $S_{X}$ and $S_{Y}$ being open intervals, we conclude that the restricted CDFs ${F_{X}}|_{S_{X}},{F_{Y}}|_{S_{Y}}$ are continuous and strictly monotonically increasing on $S_{X},S_{Y}$, which is the main observation needed in order to generalize.
The Significance of $h$ Being a Homeomorphism.
The fact that $h$ is a homeomorphism is significant in the context of generative models for several reasons:
1. Learnability – Since $h$ being a homeomorphism implies that both $h$ and $h^{-1}$ are continuous real-valued functions over some subset of $\mathbb{R}$, then by the universal approximation theorem they can be approximated to arbitrary accuracy by a neural network (Hornik, Stinchcombe, and White 1989). This is demonstrated in Figure 3 (A).
2. Coverage and diversity – The bijectivity of a homeomorphism means that there is a one-to-one and onto correspondence between the supports $S_{X}$ and $S_{Y}$. This means that we can use $Y$ to cover all samples that can be obtained from $X$ and that no two samples in $Y$ will generate the same sample in $X$. This is demonstrated in Figure 3 (A) and (B).
3. Continuous interpolation – The fact that $h$ is continuous is significant for purposes of interpolation. For example, given a generative model $g:\mathbb{R}\to S_{X}$, two points $y_{1},y_{2}\in\mathbb{R}$, and the function: $\phi(\lambda)=\lambda y_{1}+(1-\lambda)y_{2}$ (where $\phi:[0,1]\to\mathbb{R}$) we would like $(g\circ\phi)(\lambda)$ to be continuous (e.g. for video generation). If $g$ is stochastic, for example, this cannot be guaranteed. However, using $h$ as $g$, we have that $g\circ\phi$ is continuous as a composition of continuous functions. This provides the desired continuous interpolation between points from $X$. This is demonstrated in Figure 5.
4. Guiding topological properties – There are useful properties that are invariant under homeomorphisms (”topological properties”) that can guide us in designing better generative models. For example, one property is that homeomorphic manifolds must have the same dimension. The use of this property is demonstrated under Example: Swiss-Roll Parameter, where it will lead us to the conclusion that we are better off generating swiss-roll samples from a $1$-dimensional standard normal distribution, rather than from a $2$-dimensional one. Another example for a useful property is in Appendix A.3.
Figure 3: (A) Test results for a GTN $\hat{h}$ trained to map from $Y\sim\mathcal{N}(0,1)$ to the swiss-roll parameter. The color indicates which point in the normal sample was mapped to which point in the swiss-roll. (B) Test results for a GTN $\hat{h}$ trained to map from $Y\sim\mathcal{N}(0,I)$ to $X\sim U\big{(}(0,1)\times(0,1)\big{)}$. The color is based on the normal sample, i.e. for each $y$ in the normal sample, $\hat{h}(y)$ has the same color as $y$ so that the figure on the right shows how the normal sample was stretched to a uniform distribution (which is in line with Equation 4).
Learning $h$ Using a Neural Network $\hat{h}$ and Generating Samples with $\hat{h}$.
As explained in the previous subsection, one consequence of $h$ being a homeomorphism is that it can be approximated by a feedforward neural network $\hat{h}$. If we had known both $F^{-1}_{X}$ and $F_{Y}$ then we would be able to easily generate labels for each $y\in Y$: denoting the label of a given $y\in Y$ as $x_{y}$, we would set $x_{y}=h(y)$, meaning:
$x_{y}=h(y)=F_{X}^{-1}(F_{Y}(y))$ (3)
Although we do not know $F_{X}$, $F_{Y}$, we do have access to samples from $X$ and $Y$ which we can use to approximate $x_{y}$. From Eq. 3 we see that $x_{y}$ satisfies $F_{X}(x_{y})=F_{Y}(y)$, meaning $x_{y}$ is the unique $x\in X$ that matches the percentile of $y$. We can use this observation to approximate $x_{y}$ by replacing $F_{X}$ and $F_{Y}$ with the empirical CDFs as follows.
Let $D_{X}:=\{x_{1},\dots,x_{n}\}$ and $D_{Y}:=\{y_{1},\dots y_{n}\}$ be $n$ observed values of $X$ and $Y$, correspondingly. Labels are obtained easily by sorting $D_{X}$ and $D_{Y}$ to obtain $D_{X}^{\text{sorted}}$ and $D_{Y}^{\text{sorted}}$ and labeling each $y\in D_{Y}^{\text{sorted}}$ with the corresponding $x\in D_{X}^{\text{sorted}}$ (i.e. assuming $D_{X},D_{Y}$ are sorted, $y_{i}$ is assigned $x_{i}$). This is illustrated in Figure 2.
The loss function is the MSE:
$\frac{1}{n}\sum_{i=1}^{n}||\hat{h}(y_{i})-x_{y_{i}}||^{2}$
Generating samples from $X$ is now straightforward – we sample $y\in Y$ and compute $\hat{h}(y)$.
Example: Swiss-Roll Parameter.
In this example, we will see that samples from the swiss-roll can be easily generated with visibly near-perfect accuracy (Figure 3 (A)) using this method. This would not be as easy using a $2$-dimensional normal distribution (the approach taken by previous methods such as (Sohl-Dickstein et al. 2015a)), as we will now explain.
The swiss-roll is a $1$-dimensional manifold in a 2-dimensional space $\mathbb{R}^{2}$ (generated using a single, 1-dimensional, continuous random variable $\theta$). Being a $1$-dimensional manifold, it cannot be homermorphic to the support of a $2$-dimensional normal distribution ($\mathbb{R}^{2}$) since homeomorphic manifolds must have the same dimension. Therefore, there would be no hope of learning a homeomorphism that maps a 2-dimensional Gaussian random variable to the swiss-roll. However, the swiss roll is homeomorphic to $\mathbb{R}$ – the support of a $1$-dimensional standard normal distribution. In fact, since $\theta$ is a random variable that satisfies the conditions in Theorem 2.1, it is homeomorphic to $\mathbb{R}$ via $h$, which we can learn using $\hat{h}$.
To train $\hat{h}$, we created a dataset of $n=50,000$ samples from $\theta$, denoted $D_{X}$. We sampled $n$ samples from a $1$-dimensional standard normal distribution to obtain $D_{Y}$. We labeled each $y_{i}\in D_{Y}$ with its $x_{y_{i}}\in D_{X}$ as defined in the previous section. We used a standard feedforward neural network as $\hat{h}$ ($3$ layers of width $5$ with ReLU activation and batch normalization) and MSE as the loss function. Figure 3 (A) shows the result of testing the trained model $\hat{h}$ on a set of new samples $y_{1},\dots,y_{k}$ drawn from $\mathcal{N}(0,1)$ (each point is obtained by predicting $\hat{\theta}_{i}:=\hat{h}(y_{i})$ and applying the formula for the swiss roll to $\hat{\theta}_{i}$).
For comparison, we refer the reader to Figure 1 (middle) in (Sohl-Dickstein et al. 2015a) where a trained diffusion model was applied to points sampled from a standard 2-dimensional normal distribution. The points are gradually and stochastically moved back towards the swiss-roll manifold, but the result contains points that fall outside of this manifold. With $\hat{h}$, the generated points lie entirely within the data manifold. In fact, this holds true early in the training process, as shown in Appendix Figure 6.
Besides demonstrating that $\hat{h}$ can serve as an accurate generative model, this example also emphasizes an important point – namely, that if we want to learn such a homeomorphism $h$, we might need to reduce the dimensionality of the data if it is not already in its intrinsic dimension (for the swiss-roll this was $1$). This observation will guide us in later sections.
2.2 Higher Dimensions
To generalize to more than one dimension, we reduce to the $1$-dimensional case by considering lines that pass through the origin. Briefly, we take the random variables obtained by restricting $X$ and $Y$ to each line, and apply the homeomorphism defined for the $1$-dimensional case to these random variables. We begin with a formal setup that is very similar to the $1$-dimensional case.
Defining $h$.
Let $X=(X_{1},\dots,X_{d}),Y=(Y_{1},\dots,Y_{d})$ be multivariate random-variables with continuous joint probability density functions $f_{X}$, $f_{Y}$, and with supports $S_{X}$, $S_{Y}$, each of which is a product of $d$ open intervals in $\mathbb{R}$ (see Appendix A.3. to accommodate for other supports). For example, $X$ could be uniformly distributed with support $S_{X}=(0,1)\times(0,1)$ and $Y$ could be normally distributed with support $S_{Y}=\mathbb{R}\times\mathbb{R}$.
For distributions that are symmetric about the origin (e.g. standard normal), we can define the following homeomorphism (we assume for simplicity that $S_{Y},S_{X}=\mathbb{R}^{d}$):
$\begin{split}h&:\mathbb{R}^{d}\to\mathbb{R}^{d}\\ h&(y)=\begin{cases}h_{1}(||y||)\frac{y}{||y||},\quad y\neq 0\\ 0,\quad y=0\\ \end{cases}\end{split}$ (4)
where $h_{1}:(0,\infty)\to(0,\infty)$ is the $1$-dimensional homeomorphism applied to the random variables $||Y||,||X||$.
We prove that this is indeed a homeomorphism in Appendix A.5. Note that this generalizes the $1$-dimensional case since for $d=1$ we get $y\mapsto h_{1}(y)$. Intuitively, $h$ can be seen as shrinking $y$ to the unit vector in the same direction, and stretching(or shrinking it to reach the resulting $x$. This stretching and shrinking occurs on the same line through the origin, shared by both $x$ and $y$, where $h_{1}$ guarantees that the resulting $x$ has the same quantile as $y$ on this line (i.e. when measured with respect to the random variables obtained by restricting $Y,X$ to this line). Another way of thinking about this is provided in Appendix A.6.
Learning $h$.
We now turn to our labeling approach. Let $D_{X}$ and $D_{Y}$ be observed values from $X$ and $Y$. Ideally, we would use the $1$-dimensional labeling scheme on each line that passes through the origin by using $h_{1}$ on both sides of the line. In practice, however, the probability that any given line contains points from either $D_{X}$ or $D_{Y}$ is negligent. We therefore approximate this labeling approach using cosine-similarity. Specifically, we match $y$ with the $x$ that is closest in terms of both cosine-similarity (so that both $y,x_{y}$ are as close as possible to being on the same line and in the same direction from the origin) and distance from the origin (to produce the effect of $h_{1}(||y||)$). This is formally described in Algorithm 1. To see why Algorithm 1 approximates $h$, consider a line with an infinite sample from both $X$ and $Y$ in both directions from the origin. The maximal cosine-similarity over all $x\in D_{X}$ and a given $y$ on this line is $1$. Since the algorithm returns as $x_{y}$ an $x$ with maximum cosine-similarity, then $x_{y}$ will indeed be on the line and in the same direction as $y$. Since there are infinitely many such candidates $x$ with cosine-similarity $1$, it remains to choose the particular $x$ whose distance from the origin best approximates $h_{1}(||y||)$. Since both $D_{X}$ and $D_{Y}$ are sorted in ascending order by distance from the origin, and since matched pairs $(x_{y},y)$ are always removed, the algorithm will indeed return the particular $x$ from this pool of candidates whose distance from the origin will have a similar empirical quantile to that of $||y||$. We further elaborate on Algorithm 1 in Appendix A.2.
Using the labeled samples from Algorithm 1, we train $\hat{h}$ – a feedforward neural network – using MSE as the loss function. We then use $\hat{h}$ to generate new $X$ instances just as in the $1$-dimensional case and as is formally described in Algorithm 2.
Note that the requirement of respecting the $1$-dimensional homeomorphisms on each line facilitates the desired continuous interpolations of the form described in the Method section. This advantage is later demonstrated in Figure 5.
Algorithm 1 Labeling
Input:
$D_{X}=\{x_{1},\dots,x_{n}\}$
$\quad\quad\quad D_{Y}=\{y_{1},\dots,y_{n}\}\text{ sampled from }\mathcal{N}(0,I)$
Output: res
1:  $\textit{res}\leftarrow[]$
2:  $D_{X}^{\text{sorted}},D_{Y}^{\text{sorted}}\leftarrow\text{ sort }D_{X},D_{Y}\text{ ascending by $||\cdot||_{2}$}$
3:  while $D_{Y}^{\text{sorted}}\neq\emptyset$:
4:  $\quad\quad y\leftarrow D_{Y}^{\text{sorted}}[0]$
5:  $\quad\quad x_{y}\leftarrow\operatorname*{arg\,max}_{x\in D_{X}^{\text{sorted}}}\textit{cosine\_sim}(x,y)$
6:  $\quad\quad D_{Y}^{\text{sorted}}\leftarrow D_{Y}^{\text{sorted}}[1\dots n]$
7:  $\quad\quad D_{X}^{\text{sorted}}\leftarrow D_{X}^{\text{sorted}}\setminus\{x_{y}\}$
[…]
9:  return res
Algorithm 2 Sampling
Input: $\hat{h}$ trained on res (see Alg. 1)
1:  $y\leftarrow\text{ sample from }\mathcal{N}(0,I)$
2:  return $\hat{h}(y)$
Example: 2-Dimensional Uniform Distribution.
Figure 3 (B) shows a sample generated using Algorithm 2 after applying the multivariate method to the multivariate uniform distribution $X\sim U(0,1)\times U(0,1)$. Specifically, we created $D_{X}$ by sampling $n=100,000$ points from $X$, and created $D_{Y}$ by sampling the same number of points from $\mathcal{N}(0,I)$. We then applied Algorithm 1 to $D_{X},D_{Y}$ to generate labeled data. We trained a standard feedforward neural network $\hat{h}$ ($5$ hidden layers of width $6$ with ReLU activation and batch normalization (Ioffe and Szegedy 2015)) and used it to generate new samples according to Algorithm 2. The model was trained until convergence on a separately generated validation set with $n=10,000$. The colors in both images in Figure 3 (B) are based on the distance of the points from the origin in the Gaussian sample (the left image in (B)), so that the right image reflects where each point in the Gaussian was predicted to ’move’ to by $\hat{h}$.
Figure 4: Randomly chosen samples generated by GTN during the training process on MNIST at four points of improvement on the validation set. Each epoch shows 100 images, each of which is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(0,I)$). The rightmost grid contains 100 images each of which is the decoded latent vector of a random real image.
3 Experiments
In the previous section, we demonstrated how to learn $\hat{h}$ on synthetic data – the swiss-roll and the multivariate uniform distribution. In this section, we will demonstrate how to apply $\hat{h}$ to images.
In light of the discussion in Section 2.1 on using the intrinsic dimension of the data (or lower), we use autoencoders to represent each dataset in a lower dimensional latent space before training $\hat{h}$. Each dataset was trained separately (with its own autoencoder and $\hat{h}$). In all experiments, we use the same autoencoder architecture – a simple vanilla autoencoder (the encoder consisted of $2$ convolution layers with ReLU activation followed by a fully connected layer). For $\hat{h}$ we use a simple feedforward neural network with the width and depth depending on the data (as reported below) and using batch normalization and ReLU activation after each layer except for the output layer.
For $\hat{h}$, we set $X$ to be the latent vectors of the training set (the encoded images), and set $Y$ to be the standard multivariate normal distribution of the same dimension as the latent space (e.g., in CelebA we used a latent space of $100$ in the autoencoder, so $Y$ has dimension $100$). Images were generated using Algorithm 2 by computing: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ where $r\sim\mathcal{N}(0,I)$ (generation occurs in the latent space).
Figure 4 shows random sets of generated images for MNIST at several epochs of improvement on the validation set, as well as a sample of random real images for comparison. These results show that the generated samples were already realistic even after one epoch (a similar observation was also made in the swiss-roll example). We use a latent dimension of $5$ to allow for comparison with VAE for the same latent dimension (Figure 5(b) in (Kingma and Welling 2013)). Further comparisons can be made with diffusion (Figure App.1. in (Sohl-Dickstein et al. 2015b)) and a recent continuous normal flow (CNF) (Figure 3(a) (right) in (Huang and Yeh 2021)). For $\hat{h}$ we used width $50$ and depth $5$. Training was stopped when no improvement in the validation MSE was obtained for $3$ epochs.
Next we applied the method to CelebA (Figure 1 and Appendix Figure 7). Appendix Figure 7 also shows the progress in image generation during training, with randomly generated images at epochs of improvement in the Inception Score (IS) (Salimans et al. 2016). We used IS as the stopping criteria, as explained in Appendix A.4. One epoch took just under $1$ minute ($50.6$ seconds on average) on a single T4 GPU, reflecting $9$ hours until the last improvement in IS at epoch $640$. Observing Appendix Figure 7, we can see that realistic images were already obtained at half this time (see epoch $276$). Appendix Figure 8 contains a plot of the IS by epoch showing that the true IS (the IS score for a random set of real decoded samples) is nearly achieved by the GTN. The time to generate $1,000$ samples (i.e. to compute: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$) on a standard PC was $4.07$ seconds (reflecting $4.07$ milliseconds per image) and $0.00367$ seconds (reflecting $0.00367$ milliseconds per image) on a T4 GPU. We used a latent dimension of $100$ in the autoencoder. For $\hat{h}$ we used width $1200$ and depth $25$. We used a center crop of $148\times 148$ and resized the images to $64\times 64$.
Interpolation.
Figure 5: Interpolations produced by a trained GTN for CelebA (top) and the HaP dataset (bottom).
We next used the CelebA dataset as well as the Hands and Palm Images (HaP) dataset (Kaggle. 2024) to demonstrate that GTNs indeed provide continuous interpolations. The HaP dataset is a small dataset containing $11,000$ images of upward-facing and downward-facing hands. We chose the HaP dataset since generating hands is known to be notoriously difficult for generative models (The New Yorker. 2023; Vox. 2023), let alone interpolating between them.
Figure 5 shows that GTN indeed generates continuous interpolations for both CelebA and HaP. Each grid contains generated interpolations between two generated images (top left and bottom right in each grid). More precisely, each grid contains the results of decoding $\hat{h}\big{(}\lambda y_{\text{top\_left}}+(1-\lambda)y_{\text{bottom\_right}}\big{)}$ for $100$ linearly spaced $\lambda\in[0,1]$ where both $y_{\text{top\_left}}$ and $y_{\text{bottom\_right}}$ were sampled from $\mathcal{N}(0,I)$.
It is worth noting an interesting observation for the HaP data results – the interpolations in Figure 5 are between hands of the same orientation (downward facing). However, generated interpolations between hands (downward facing) and palms (upward-facing) do not seem as natural (Appendix Figure 9). This is likely because the dataset does not contain in-between positions for the latter (transitions between downward and upward), but it does contain in-between positions for the former (closed/open fingers to various degrees). This demonstrates that, regardless of methodology, the completeness of the dataset is important for accurate image generation, and particularly for interpolation.
4 Related work
GTNs are related in concept to NFs which seek to map a source distribution into a target distribution using a sequence of bijective transformations. These are typically implemented by a neural network, often in the form of at least tens of neural network blocks (Xu and Campbell 2024) and sometimes more (Xu, Chen, and Campbell 2023). The core of NFs is based on each of these transformations being a diffeomorphism – a notion similar to a homeomorphism but more constrained as it is defined between smooth manifolds (as opposed to any topological spaces, including any manifolds) and requires that the function and its inverse are differentiable (as opposed to just continuous for homeomorphisms). NFs also require that the log-determinant of the Jacobian of these transformations is tractable, posing a constraint on the model architecture. GTNs do not pose any limitations on the model architecture. NFs also differ from GTNs in their optimization method since NFs employ a maximum likelihood objective – typically the Kullback-Leibler (KL) divergence while GTNs are trained using MSE. It was observed that optimizing the KL-divergence may be difficult for more complex normalizing flow structures (Xu, Chen, and Campbell 2023). NFs may also suffer from low-quality sample generation (Behrmann et al. 2021), partially because the constraint on the Jacobian, besides limiting the model structure, may also lead to issues such as spurious local minima and numerical instability (Golinski, Lezcano-Casado, and Rainforth 2019).
Continuous normalizing flow (CNF) are a relatively recent type of NFs that were developed to avoid the main constraints posed by NFs – namely the requirements of invertibility and a tractable Jacobian determinant. To avoid these constraints, CNFs use ODEs to describe the transformations in NFs. While CNFs have shown improvements over NFs in certain cases , they are extremely time-intensive since they require solving a series of ODEs, equivalent to running a neural network with hundreds of layers (Huang and Yeh 2021). Improving both the speed and performance of CNFs is an active field of research. To compare between a recently suggested CNF and GTN see the Experiments section.
5 Discussion
In this work we introduced a new class of generative models – GTNs – that learn a homeomorphism $h$ between the supports of two continuous random variables using samples from both. The samples are used in an efficient algorithm to create a training dataset for the GTN which is trained in a simple supervised approach, with no restrictions on the model architecture. The freedom to use a standard neural network architecture, and the fact that $h$ satisfies the conditions of the universal approximation theorem, lead GTNs to quickly converge towards $h$, providing an efficient method to generate samples from a target random variable. This was demonstrated experimentally in multiple cases, particularly in that generated samples reached the data manifold at early epochs and in the total time to convergence. Other positive implications of $h$ being a homeomorphism – namely, coverage and diversity, continuous interpolations and guiding topological properties – were also discussed and demonstrated experimentally.
One downside of our method is that it assumes that the data is in its intrinsic dimension (or lower). This may require use of dimensionality reduction methods, such as the vanilla autoencoders employed in this work. Future work focused on high resolution images may be interested in experimenting with stronger autoencoders. Additionally, the assumptions on the support of the distributions mean that our method does not immediately apply to discrete distributions, and that datasets may need to be split and trained on individually, as explained in Appendix A.3. This also means that larger datasets may be required for GTNs to perform well.
6 Broader Impact
Our work falls under the category of deep generative models, and is similar in impact to other studies seeking to improve these methods. Since our work provides an efficient method to generate realistic samples, it may contribute to malicious uses such as generating fake news, videos of high profile figures etc. Fortunately, there are methods to detect such generated data, and there is ongoing work to continue to improve in doing so (Wang et al. 2020). There are also concerns of perpetuating biases inherent to the dataset, particularly since our method essentially learns the data distribution with high fidelity. There is a growing line of research on detecting and handling biases in datasets, and we believe there is an increasing awareness among data curators and users of this problem. More research would be necessary in order to alleviate this problem.
Although malicious uses are indeed a concern, generative models could have far reaching positive implications on our lives. Perhaps the most prominent example is the field of healthcare, where generative models have found use in generating digital twins for improved personalized medicine (Bordukova et al. 2024; Tsialiamanis et al. 2021), medical imaging (Kazerouni et al. 2023; Moghadam et al. 2023; Pinaya et al. 2022) and drug design (Korablyov et al. 2024; Chen et al. 2023; Hoogeboom et al. 2022; Schneuing et al. 2022; Corso et al. 2022).
Appendix A Appendix
A.1
  1. 1.
    $h$ is well-defined: we need to show that: (a) ${F_{X}}|_{S_{X}}$ is invertible, and (b) that the image of ${F_{Y}}|_{S_{Y}}$ is in the domain of ${F_{X}}|_{S_{X}}$. For (a), observe that ${F_{X}}|_{S_{X}}$ is continuous, strictly monotonically increasing (the CDF is strictly monotonically increasing on the support) and onto $(0,1)$. A symmetric argument can be made for ${F_{Y}}|_{S_{Y}}$, which we will use later to define the inverse of $h$.
    For (b), we know from (a) that the image of ${F_{Y}}|_{S_{Y}}$ is the domain of ${F_{X}}|_{S_{X}}$ so $h$ is well defined.
  2. 2.
    $h$ is a homeomorphism since: a. it is continuous as a composition of continuous functions (${F_{X}^{-1}}|_{S_{X}}$ is the inverse of a continuous and strictly monotonous function on an open interval in $\mathbb{R}$); b. it is bijective with inverse ${F_{Y}^{-1}}|_{S_{Y}}({F_{X}}|_{S_{X}}(x))$; c. its inverse is continuous for the same reasons stated in a. Also, it is a known result that a continuous and bijective function between two open intervals has a continuous inverse.
A.2
To understand the rationale behind Algorithm 1, imagine four points from each of $X$ and $Y$ lying on the $x$-axis in $\mathbb{R}^{2}$ such that both $X$ and $Y$ have two points on each side of the origin (we assume $D_{X},D_{Y}$ are centralized). The cosine similarity of each pair of points $\{a,b\}$ with $a\in D_{X},b\in D_{Y}$ is $1$ if they are on the same side or $-1$ of they are on opposite sides. Take the first $y\in D_{Y}^{\text{sorted}}$ (the closest one to the origin which reflects $\sim 50$th percentile since we assumed an equal number of points from $Y$ on both sides). The maximum cosine similarity out of all $x\in D_{X}^{\text{sorted}}$ is $1$, meaning $y$ will be matched with an $x$ on the same side. The fact that $D_{X}$ is sorted, means that we are using the distance as a tie breaker – out of all $x\in D_{X}$ that are on the same side as $y$, the $x$ that is closest to the origin will be chosen ($\sim 50$th percentile). Likewise, the leftmost $y$ will be matched with the leftmost $x$ (both 00th percentile), the rightmost $y$ with the rightmost $x$ (both $100$th percentile) etc.
A.3
One example of a topological property is the connectedness property – being a ”single piece”, like $\mathbb{R}$, as opposed to ”more than one piece”, like $(-\infty,-1)\cup(1,\infty)$, is preserved between homeomorphic spaces (so, in particular, these two are not homeomorphic, but $\mathbb{R}$ is homeomorphic to each of the two intervals, separately). Therefore, if we suspect that our data is composed of separate classes like images of both hands and faces, and we wish $\hat{h}$ as a generative model, we could split our data to its different components (learning a separate $h$ for each). This relates to the conditions of Theorem 2.1 which assume that $X$’s support is an open interval. For example, if we suspect that our data can be separated into two disjoint uniform distributions, then the theorem doesn’t apply. However, it does apply to each separate uniformly distributed component.
Figure 6: Swiss-roll samples generated by GTN during the training process.
Figure 7: Randomly chosen samples generated by GTN during the training process on celebA at several epochs of improvement in the IS score, and randomly chosen real images (bottom-right grid). Each epoch shows 200 images, each of which is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(0,I)$). The bottom right grid contains 200 images each of which is the decoded latent vector of a random real image.
A.4
We experimented with two stopping criteria. The first was based on the validation set, as in the MNIST case. The second used IS to assess the quality of the generated images using a small random set of $200$ images. IS yielded better results. Training was stopped when no improvement in the IS score was obtained for $300$ epochs, but observing the generated samples by epoch above demonstrates that earlier stopping may suffice.
Figure 8: Inception score (IS) for $200$ randomly generated images (decoded latent vectors generated by $\hat{h}$ from random normal vectors) plotted by epoch (at epochs of improvement in the score) (blue); and IS of a random set of $200$ real images (decoded latent vectors of real images) (green).
Figure 9: Interpolations between a hand (downward facing) and a palm (upward facing) (left) and vice versa (right).
A.5
$h$ is injective as can be observed by taking the norm when equating for two different points $y_{1},y_{2}$:
$h_{1}(||y_{1}||)\frac{y_{1}}{||y_{1}||}=h_{1}(||y_{2}||)\frac{y_{2}}{||y_{2}||}$
which yields $h_{1}(||y_{1}||)=h_{1}(||y_{2}||)$. Since $h_{1}$ is injective, this means that $||y_{1}||=||y_{2}||$. Plugging this into the equality and cancelling out equal terms ($y_{1},y_{2}\neq 0)$ yields $y_{1}=y_{2}$. The function is also onto since $h_{1}$ is onto. Particularly, for any $x\neq 0$ there is a value $v$ in $(0,\infty)$ with $h_{1}(v)=||x||$. Since the support of $Y$ is $\mathbb{R}^{d}$ there is a $y$ that is on the same line as $x$ from the origin and that satisfies $||y||=v$ so that: $x=h_{1}(||y||)\frac{y}{||y||}$.
$h$ is continuous since for $y\neq 0$ it is a composition of continuous functions, and for $y=0$ we can observe that $\lim_{y\to 0}h_{1}(||y||)\frac{y}{||y||}=0$ ($\frac{y}{||y||}$ is the unit vector in the direction of $y$ and $h_{1}(||y||)\to 0$ as $y\to 0$ by definition of the CDF).
The inverse is also continuous since a continuous and bijective functions between open subsets of $\mathbb{R}^{d}$ is continuous (by the invariance of domain theorem).
A.6
Another way of thinking about this is as follows: assume for simplicity that $f_{X},f_{Y}>0$ (so that $S_{X},S_{Y}=\mathbb{R}^{d}$ and therefore $l\cap S_{X}$ and $l\cap S_{Y}$ are simply $l$). Since $l$ is a $1$-dimensional manifold, $X$ and $Y$ induce $1$-dimensional random variables $X^{l}$ and $Y^{l}$ on the line $l$, with pdfs $f_{X^{l}},f_{Y^{l}}$ and CDFs $F_{X^{l}},F_{Y^{l}}$. We can apply Theorem 2.1 to these random variables to obtain a homeomorphism $h^{l}$ for each line. As mentioned earlier, under certain conditions on $X,Y$, this process yields a multivariate homeomorphism.